# Fast and accurate parameter estimation of high-redshift sources with the Einstein Telescope

### v1

@@ V Conclusions
The Einstein Telescope (ET), along with other third-generation (3G) detectors such as Cosmic Explorer (Reitze et al., 2019; Evans et al., 2023; Gupta et al., 2024) and next-generation detectors like LISA (Colpi et al., 2024), will define the future of gravitational wave (GW) astronomy in the coming decades (Punturo et al., 2010; Maggiore et al., 2020; Branchesi et al., 2023). These detectors are expected to observe thousands of sources every day (Borhanian and Sathyaprakash, 2024). However, this enormous volume of data will pose significant challenges for data analysis and parameter estimation (PE; Couvares et al., 2021). Current inference algorithms, which rely on stochastic sampling (Ashton et al., 2019), will struggle to handle this scale of data efficiently and rapidly, even with techniques that implement likelihood simplification (Narola et al., 2024; Hu and Veitch, 2024) or advanced sampling methods (Williams et al., 2023; Wong et al., 2023a). Therefore, the era of 3G detectors requires new technologies for PE, which must be developed now to guide the design of these detectors.
@@
To tackle this challenge, we focus on providing fast and accurate PE for a specific category of sources which is pivotal in the science case of 3G detectors: high-redshift BBH mergers (Franciolini et al., 2022; Ng et al., 2021, 2022a, 2022b; Santoliquido et al., 2023, 2024; Reali et al., 2024). These short-lived sources are well-suited for being processed with Neural Posterior Estimation (NPE; Papamakarios and Murray, 2016), a likelihood-free inference method implemented in Dingo-IS (Green and Gair, 2021; Dax et al., 2021, 2025). NPE learns the posterior probability distribution directly from simulated data using normalizing flows (Kobyzev et al., 2021).
@@
We demonstrate in this work that NPE is a promising approach for 3G detector applications. With sufficient training and samples refined through importance sampling (see Sec. II.4), the posterior distributions obtained from NPE are statistically indistinguishable from those derived using standard inference methods (Sec. III.1 and Fig. 5), at a fraction of time and energy cost (Sec. IV.5).
@@
The triangular configuration of ET (ET-$\Delta$, Fig. 1) adopted in this work revealed unique characteristics in the posterior probabilities of detected sources, particularly in sky localization. We show that ET-$\Delta$ will observe up to eight degenerate sky modes for short-lived distant sources (Sec. IV.1 and Fig. 10), which can be explained by time- and frequency-independent antenna patterns.
@@
We analyze $1000$ simulated BBH signals extracted from the prior distribution (Sec. III.2) and achieve high accuracy in parameter estimation across a wide redshift range ($1\lesssim z\lesssim 45$). The sampling efficiency is generally higher for more distant BBHs, averaging 12.5% overall and increasing to about 20% for sources merging at $z>10$ (Fig. 7). This trend is confirmed in an analysis of 690 sources from a one-year observation of BBHs formed from Population III stars (Costa et al., 2023; Santoliquido et al., 2023, 2024), where the average sampling efficiency for $z>10$ is about 16% (Sec. IV.3 and Fig. 13).
@@
This study compares parameter estimation accuracy obtained with NPE and with the FIM formalism for the triangular ET configuration (Sec. III.2). On one hand, NPE and FIM show good agreement for chirp mass and mass ratio (Figs. 8 and 9), confirming previous results (Dupletsa et al., 2025). On the other hand, FIM tends to overestimate luminosity distance errors for sources merging at $z>10$, with discrepancies reaching up to a factor of 4. In sky localization, FIM fails to capture the eight degenerate sky modes characteristic of a triangular ET, significantly underestimating errors.
@@
These findings highlight the limitations of FIM in accurately capturing complex parameter dependencies as they can emerge in observations with a single ET triangle. However, as discussed in Sec. IV.4, when time- and frequency-dependent antenna patterns are taken into account (Singh and Bulik, 2022; Piro et al., 2023), FIM analysis is expected to provide sky localizations comparable to those obtained with full PE. This holds for BNS mergers, supporting the validity of previous FIM-based studies (Borhanian and Sathyaprakash, 2024; Iacovelli et al., 2022; Ronchini et al., 2022; Banerjee et al., 2023; Branchesi et al., 2023; Loffredo et al., 2024).
@@ Acknowledgements.
We thank Luca Reali, Francesco Iacovelli, Andrea Maselli, Lavinia Paiella, Cristiano Ugolini, Benedetta Mestichelli, Alessio Ludovico De Santis, and Matteo Schulz for insightful discussions. F.S. acknowledges financial support from the AHEAD2020 project (grant agreement n. 871158). M.B. acknowledges financial support from the Italian Ministry of University and Research (MUR) for the PRIN grant METE under contract no. 2020KB33TP. MAS acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101025436 (project GRACE-BH, PI: Manuel Arca Sedda). MAS acknowledge financial support from the MERAC foundation. S.R.G. was supported by a UK Research and Innovation (UKRI) Future Leaders Fellowship (grant no. MR/Y018060/1). I.M.R.-S. acknowledges the support of the Herchel Smith fund. E.B. is supported by NSF Grants No. AST-2307146, No. PHY-2207502, No. PHY-090003 and No. PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant No. 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. The research leading to these results has been conceived and developed within the Einstein Telescope Observational Science Board ET-0130A-25.
@@ Appendix A Priors
Fig. 14 shows the distribution of injected parameters sampled accordingly to the chosen priors as presented in Sec. II.2 and listed in Tab. 1. We show only the non-trivial parameters like detector-frame chirp mass, mass ratio, luminosity distance, and aligned spins, since the other parameters are distributed uniformly in the specified ranges. The primary (secondary) detector-frame mass is constrained to be within 50 M and 1500 M (1000 M). This explains the sharp edges in the $p(\mathcal{M}_{d},q)$ distribution in Fig. 14.
Figure 14: Marginalized one- and two-dimensional distributions of the injection parameters, sampled according to the chosen priors listed in Tab. 1. Contours indicate the 68%, 95%, and 99% credible regions. For details, see Sec. II.2 and Appendix A.
@@ Appendix B Event 2
Figure 15: Marginalized one- and two-dimensional posterior distributions for Event 2 over all set of parameters, comparing Dingo-IS (light blue) and Bilby (orange). Vertical and horizontal lines mark the true injected values (see Tab. 4). Contours represent 68% and 95% credible regions. Sky modes are discussed in Sec. IV.1. See Appendix B for other details.
For completeness, we present the same analysis as in Sec. III.1 but for a different injection, Event 2, which has a higher injected redshift and therefore a lower SNR with respect to Event 1. We draw $1\times 10^{5}$ samples from $q(\theta|d)$. The sampling efficiency for Event 2 is approximately $25\%$, corresponding to $\sim 2.5\times 10^{4}$ effective samples. In Bilby we set the number of live points to nlive = 10_000, which takes $>20$ hours to get to convergence.
As shown in Tab. 4, the JSD values for the one-dimensional marginal distribution of all parameters except the chirp mass ($\mathcal{M}_{\mathrm{d}}$), inclination angle ($\theta_{\mathrm{jn}}$), and phase ($\phi_{c}$) are all $<2\times 10^{-3}$ nat, indicating that the posteriors obtained with Dingo-IS and Bilby are identical. The remaining parameters exhibit small but non-negligible differences. This is also evident in Fig. 15, which displays the one- and two-dimensional marginal posteriors. In particular, Dingo-IS produces marginal distributions with longer tails, which is a characteristic of NPE (Dax et al., 2023).
The slight difference between the posterior distributions obtained with the two methods is further emphasized by the difference in their computed Bayesian evidence values. The log-evidence estimates are $\log\mathcal{Z}(d)=-5917.651\pm 0.005$ for Dingo-IS and $-5917.90\pm 0.04$ for Bilby, a statistically significant difference.
To get a difference between the two posteriors which is statistically insignificant, as in the case of Event 1 in Sec. III.1, it is necessary to increase the number of posterior samples from the two methods, which in the case of Bilby translates in increasing the number of live points and consequently the sampling time.
Parameter $\theta^{\mathrm{inj}}$ JSD×$\times 10^{3}$ $\theta$
$\mathcal{M}_{\mathrm{d}}$ 909 M 2.5 nat ${805}_{-103}^{+80}$ M
$q$ 0.89 2.0 nat ${0.8}_{-0.3}^{+0.2}$
$d_{\mathrm{L}}$ 413 Gpc 0.9 nat ${153}_{-83}^{+108}$ Gpc
ra 3.94 rad 0.8 nat ${3.20}_{-2.82}^{+2.35}$ rad
dec 0.97 rad 1.7 nat ${-0.00}_{-1.03}^{+1.04}$ rad
$\theta_{\mathrm{jn}}$ 2.82 rad 3.4 nat ${1.73}_{-0.97}^{+0.72}$ rad
$\phi_{c}$ 3.21 rad 2.6 nat ${4.22}_{-3.17}^{+1.22}$ rad
$\psi$ 0.13 rad 1.2 nat ${1.60}_{-1.36}^{+1.35}$ rad
$t_{\mathrm{geocent}}$0.04 s 0.9 nat ${0.04}_{-0.03}^{+0.04}$ s
$\chi_{1}$ 0.42 1.9 nat ${0.01}_{-0.34}^{+0.41}$
$\chi_{2}$ 0.31 1.0 nat ${0.04}_{-0.43}^{+0.54}$
@@
Table 4: Injected parameters $\theta^{\mathrm{inj}}$ for Event 2 and Jensen-Shannon divergence (JSD) quantifying the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. The last column on the right shows the median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS. We report in the bottom rows the source-frame primary and secondary mass, redshift and optimal signal-to-noise ratio ($\rho$, see Eq. 17). See Appendix B for details.

### v2

@@ V Conclusions
The Einstein Telescope, along with other third-generation detectors such as Cosmic Explorer (Reitze:2019iox; Evans:2023euw; Gupta:2023lga) and next-generation detectors like LISA (LISA:2024hlh), will define the future of gravitational wave astronomy in the coming decades (Punturo:2010zz; ET:2019dnz; Branchesi:2023mws). These detectors are expected to observe thousands of sources every day (2024PhRvD.110h3040B). However, this enormous volume of data will pose significant challenges for data analysis and parameter estimation (PE; Couvares:2021ajn). Current inference algorithms, which rely on stochastic sampling (Ashton:2018jfp), will struggle to handle this scale of data efficiently and rapidly, even with techniques that implement likelihood simplification (2024PhRvD.110h4085N; 2024arXiv241202651H) or advanced sampling methods (2023MLS&T...4c5011W; 2023ApJ...958..129W). Therefore, the era of 3G detectors requires new technologies for PE, which must be developed now to guide the design of these detectors.
@@
To tackle this challenge, we focus on providing fast and accurate PE for a specific category of sources which is pivotal in the science case of 3G detectors: high-redshift BBH mergers (Franciolini:2022tfm; Ng:2020qpk; Ng:2021sqn; Ng:2022agi; santoliquido2023; Santoliquido:2024oqs; Reali:2024hqf). These short-lived sources are well-suited for being processed with Neural Posterior Estimation (NPE; Papamakarios:2016ctj), a likelihood-free inference method implemented in Dingo-IS (Green:2020dnx; Dax:2021myb; Dax:2024mcn). NPE learns the posterior probability distribution directly from simulated data using normalizing flows (Kobyzev_2021).
@@
We demonstrate in this work that NPE is a promising approach for 3G detector applications. With sufficient training and samples refined through importance sampling (see Sec. II.4), the posterior distributions obtained from NPE are statistically indistinguishable from those derived using standard inference methods (Sec. III.1 and Fig. 5), at a fraction of time and energy cost (Sec. IV.6).
@@
The triangular configuration of ET (ET-$\Delta$, Fig. 1) adopted in this work reveals unique characteristics in the posterior probabilities of detected sources, particularly in sky localization. We show that ET-$\Delta$ will observe up to eight degenerate sky modes for short-lived distant sources (Sec. IV.1 and Fig. 10), which can be explained by time- and frequency-independent antenna patterns.
@@
We analyze $1000$ simulated BBH signals extracted from the prior distribution (Sec. III.2) and achieve high accuracy in parameter estimation across a wide redshift range ($1\lesssim z\lesssim 45$). The sampling efficiency is generally higher for more distant BBHs, averaging 12.5% overall and increasing to about 20% for sources merging at $z>10$ (Fig. 7). This trend is confirmed in an analysis of 690 sources from a one-year observation of BBHs formed from Population III stars (costa2023; santoliquido2023; Santoliquido:2024oqs), where the average sampling efficiency for $z>10$ is about 16% (Sec. IV.4 and Fig. 14).
@@
This study compares parameter estimation accuracy obtained with NPE and with the FIM formalism for the triangular ET configuration (Sec. III.2). FIM tends to overestimate luminosity distance errors for sources merging at $z>10$, with discrepancies reaching up to a factor of 4. In sky localization, FIM fails to capture the eight degenerate sky modes characteristic of a triangular ET, significantly underestimating errors. For other parameters, such as the detector-frame chirp mass, mass ratio, and aligned spins, the FIM generally underestimates the uncertainties. These results pertain specifically to the science case investigated in this work, namely high-redshift BBH mergers.
@@
These findings highlight the limitations of FIM in accurately capturing complex parameter dependencies as they can emerge in observations with a single ET triangle. However, as discussed in Sec. IV.5, when time- and frequency-dependent antenna patterns are taken into account (Singh:2021bwn; Piro:2022zos), FIM analysis is expected to provide sky localizations comparable to those obtained with full PE (Chan:2018csa; Nitz:2021pbr; Li:2021mbo; Li:2024ruk; Singh:2021bwn).
Data Availability
The main data used in this study are publicly available on Zenodo (santoliquido_2025_17358571) (https://zenodo.org/records/17358571). The latest public release of Dingo can be accessed from https://github.com/dingo-gw/dingo, while GWFish+Priors is also openly available at https://github.com/janosch314/GWFish. Additional data and code are available from the corresponding authors upon reasonable request.
@@ Acknowledgements.
We thank the anonymous referee for their careful reading of our manuscript and for the valuable suggestions that helped improve it. We thank Michael Williams, Luca Reali, Francesco Iacovelli, Andrea Maselli, Lavinia Paiella, Cristiano Ugolini, Benedetta Mestichelli, Alessio Ludovico De Santis, and Matteo Schulz for insightful discussions. F.S. acknowledges financial support from the AHEAD2020 project (grant agreement no. 871158). M.B. acknowledges financial support from the Italian Ministry of University and Research (MUR) for PRIN grant METE under contract no. 2020KB33TP. MAS acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 101025436 (project GRACE-BH, PI: Manuel Arca Sedda). MAS acknowledge financial support from the MERAC foundation. S.R.G. was supported by a UK Research and Innovation (UKRI) Future Leaders Fellowship (grant no. MR/Y018060/1). I.M.R.-S. acknowledges the support of the Herchel Smith fund. E.B. is supported by NSF grants no. AST-2307146, no. PHY-2207502, no. PHY-090003 and no. PHY-20043; by NASA grant no. 21-ATP21-0010; by the John Templeton Foundation grant no. 62840; by the Simons Foundation; and by the Italian Ministry of Foreign Affairs and International Cooperation grant no. PGR01167. The research leading to these results has been conceived and developed within the Einstein Telescope Observational Science Board ET-0130A-25.
@@ Appendix A Priors
Fig. 15 shows the distribution of injected parameters sampled according to the chosen priors as presented in Sec. II.2 and listed in Table 1. We show only the nontrivial parameters like detector-frame chirp mass, mass ratio, luminosity distance, and aligned spins, since the other parameters are distributed uniformly in the specified ranges. The primary (secondary) detector-frame mass is constrained to be within 50 M and 1500 M (1000 M). This explains the sharp edges in the $p(\mathcal{M}_{d},q)$ distribution in Fig. 15.
Figure 15: Marginalized one- and two-dimensional distributions of the injection parameters, sampled according to the chosen priors listed in Table 1. Contours indicate the 68%, 95%, and 99% credible regions. For details, see Sec. II.2 and Appendix A.
@@ Appendix B Event 2
Figure 16: Marginalized one- and two-dimensional posterior distributions for Event 2 over all set of parameters, comparing Dingo-IS (light blue) and Bilby (orange). Vertical and horizontal lines mark the true injected values (see Table 4). Contours represent 68% and 95% credible regions. Sky modes are discussed in Sec. IV.1. See Appendix B for other details.
For completeness, we present the same analysis as in Sec. III.1 but for a different injection, Event 2, which has a higher injected redshift and therefore a lower SNR with respect to Event 1. We draw $1\times 10^{5}$ samples from the proposal distribution $q(\theta|d)$. The sampling efficiency for Event 2 is approximately $25\%$, corresponding to $\sim 2.5\times 10^{4}$ effective samples.
In Bilby, we set the number of live points to nlive = 10_000. The default settings in nessai do not yield satisfactory results for this source. To improve robustness, we increase the volume_fraction from 0.95 to 0.98. This parameter determines the fraction of the total probability enclosed by the contour in latent space—the base space from which samples are initially drawn before being mapped to physical space via the normalizing flow (2021PhRvD.103j3006W). While a higher volume_fraction enhances robustness, it also increases computational cost. With these settings, Bilby requires approximately 1 day and 20 hours to reach convergence.
We follow the same procedure outlined in Sec. III.1 to define a threshold on the JSD for assessing statistical significance. In this case, we generate with Dingo-IS 100 sets of $10^{5}$ posterior samples each.
As shown in Table 4, the median JSD values for the one-dimensional marginal distribution of all parameters are below the corresponding JSDthr, indicating that the posteriors obtained with Dingo-IS and Bilby are identical. This agreement is also evident in Fig. 16, which displays the one- and two-dimensional marginal posteriors.
The log-evidence estimates are $\log\mathcal{Z}(d)=-5917.651\pm 0.005$ for Dingo-IS and $5918.13\pm 0.04$ for Bilby, a difference within statistical uncertainty.
Parameter $\theta^{\mathrm{inj}}$ $\langle\mathrm{JSD}\rangle$ JSDthr $\theta$
$\mathcal{M}_{\mathrm{d}}$ 909 M 1.3 2.3 ${805}_{-103}^{+80}$ M
$q$ 0.89 1.1 2.3 ${0.8}_{-0.3}^{+0.2}$
$d_{\mathrm{L}}$ 413 Gpc 1.6 2.6 ${153}_{-83}^{+108}$ Gpc
ra 3.94 rad 1.0 2.7 ${3.20}_{-2.82}^{+2.35}$ rad
dec 0.97 rad 1.5 3.0 ${-0.00}_{-1.03}^{+1.04}$ rad
$\theta_{\mathrm{jn}}$ 2.82 rad 2.2 2.7 ${1.73}_{-0.97}^{+0.72}$ rad
$\phi_{c}$ 3.21 rad 1.6 2.9 ${4.22}_{-3.17}^{+1.22}$ rad
$\psi$ 0.13 rad 1.2 3.0 ${1.60}_{-1.36}^{+1.35}$ rad
$t_{\mathrm{geocent}}$ -0.04 s 1.0 2.8 ${0.04}_{-0.03}^{+0.04}$ s
$\chi_{1}$ 0.42 1.3 2.8 ${0.01}_{-0.34}^{+0.41}$
$\chi_{2}$ 0.31 1.1 2.5 ${0.04}_{-0.43}^{+0.54}$
@@
$\Delta\Omega_{90\%}$ 22 922 deg2
Table 4: Injected parameters $\theta^{\mathrm{inj}}$ for Event 2, along with the median Jensen-Shannon divergence ($\langle\mathrm{JSD}\rangle$, in units of $10^{-3}$ nat), which quantifies the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. For comparison, the corresponding JSD threshold (JSDthr) is also shown. The last column on the right shows the median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS. We report in the bottom rows the source-frame primary and secondary mass, redshift, the optimal signal-to-noise ratio ($\rho$, see Eq. 17), and the sky localization error. See Appendix B for details.